Session: Nov 2000 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Wed Nov 22 10:07:08 MET 2000) Success

Torun (K.Borkowski at Wed Nov 22 10:36:33 MET 2000) Success
No known problems.

Westerbork (Tony Foley at Wed Nov 22 11:20:40 MET 2000) Minor failures
Scans 326-1615, 327-0212, 327-0515 replaced by local calibration as they were outside hour angle range. Total power on 3C273 seems to give overflow in BBC1.

Effelsberg (A. Kraus at Thu Nov 23 12:46:09 MET 2000) Minor failures
The antenna had to be stopped because of strong wind between UT2030-2116 and UT 2200-0000.

Onsala (A. Polatidis at Fri Nov 24 16:01:13 MET 2000) Success
RFI in the band, hopefully did not affect the observed bandwidth that much

Noto (C.Trigilio at Tue Nov 28 18:13:06 MET 2000) Minor failures
Lost from 22:37 to 22:48 for cable wrap problem (antenna in prelimit).

Jodrell_Bank (Alastair Gunn at Mon Dec 4 13:21:51 MET 2000) Minor failures
Telescope did not move to 3C273 at 05:30. FS reported that telescope was on source so problem was not noticed immediately. On source from about 08:00. No other known problems.

Cambridge (Alastair Gunn at Mon Dec 4 13:22:12 MET 2000) Success
No known problems.